OpenStudy (anonymous):

While a planet P rotates in a circle about its sun, a moon M rotates in a circle about the planet, and both motions are in a plane. Let's call the distance between M and P one. Suppose the distance of P from the sun is 3.5×103 lunar units; the planet makes one revolution about the sun every 9 years, and the moon makes one rotation about the planet every 0.25 years. Choosing coordinates centered at the sun, so that, at time t=0 the planet is at (3.5×103, 0), and the moon is at (3.5×103, 1), then the location of the moon at time t, where t is measured in years, is (x(t),y(t)), x(t)=??? y(t)=???

OpenStudy (anonymous):

Fun! P's position is given by: \[P _{x}(t)=(3.5\times103)\cos (2\pi t/9)\] \[P _{y}(t)=(3.5\times103)\sin (2\pi t/9)\] We need to add the moon's position in a similar fashion; need more direction?

OpenStudy (anonymous):

says both r wrong

OpenStudy (anonymous):

That's just the planet's position. You need to add the moon's; although I'm guessing the 3.5x103 in the OP should read 3.5x10^3.

OpenStudy (anonymous):

To add the moon you also have sinusoidal functions with a shorter period and a phase shift of pi/2.

OpenStudy (anonymous):

still said they r wrong